Find an equation of the line that passes through the given point and is perpendicular to the given line. Write the equation in slope-intercept form.
(8, −4), 5x = 7y −

Found 2 solutions by mananth, 9948007849:
Answer by mananth(16946)

5 x + -7 y = -6
Find the slope of this line

-7 y = -5 x -6
Divide by -7
y = 5/ 7 x + 6/7
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 5/7

The slope of a line perpendicular to the above line will be the negative reciprocal -1 2/5
Because m1*m2 =-1
The slope of the required line will be -1 2/5

m= -7/5 ,point ( 8 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = -56/ 5 + b
b= 365
m= -7/5
The required equation is y = -7/ 5 x + 36/ 5
m.ananth@hotmail.ca

Answer by 9948007849(1)

7y=5x+6
y=5/7x+6/7
the slop line which is perpendicular to given line is -7/5
the eqation of the line which is passing throw(8,-4)and having slope -7 is
y-(-4)=-7/5(x-8)
5(y+4)=-7x+56
7x+5y-52=0